this is a differential equation question


please solve them with steps and details also, make sure it is correct

3. (6 points) Solve the IVP: y +9y = e3+ + 8(t - 27), y(0) = y0) = 0.
Answer #1

-27 te P-3 Y+gy = est + s (A-27) Applying Laplace transform to the both sides of the given equation we get _{yu} + L39Y) = L

1• ਲਾਡ - ( ਸਰਦੀਆਂ I det fit) = [{ FO} = L{ pot zn} sinat o ay to est I cost 18 6 13 sinst tu (A) f(t-21) ... L{ pla} = eat

3t » Y= casst-singt) + Usa (t). sinst 18 3 where Uso (t)=co, t <2a t 7,2A -